A container with a volume of 14 L filled with gas at 2.00 atm at 283k. What will be the new pressure temperature increase to 321k? 




1
Expert's answer
2022-03-29T11:12:06-0400

P1 = 2.00 atm

T1 = 283 K

V1 = V2 = 14 L = constant

P2 = ???

T2 = 321 K


Solution:

Since the volume and amount of gas remain unchanged, Gay-Lussac's law can be used.

Gay-Lussac's law can be expressed as: P1/T1 = P2/T2

P1T2 = P2T1

To find the new pressure, solve the equation for P2:

P2 = P1T2/T1

P2 = (2.00 atm × 321 K) / (283 K) = 2.27 atm

P2 = 2.27 atm


Answer: The new pressure will be 2.27 atm